40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 4 19:31:17 2011, kiwamu, Update, SUS, PRM and BS oplev laser died oplevs.png
    Reply  Wed Oct 5 11:33:25 2011, steve, Update, SUS, PRM and BS oplev laser replaced 
       Reply  Mon Apr 20 16:42:48 2015, steve, Update, SUS, PRM and BS oplev laser replaced newOplevLaser.png
          Reply  Fri Apr 24 23:40:15 2015, rana, Update, SUS, PRM and BS oplev laser replaced BSOL.pdf
             Reply  Sun Apr 26 15:34:34 2015, Jenne, Update, SUS, PRM oplev centered 
             Reply  Tue Jun 9 10:57:12 2015, Steve, Update, SUS, BS oplev laser tuned 100dBStrend0425.png18dBSdrift.png
Message ID: 5618     Entry time: Tue Oct 4 19:31:17 2011     Reply to this: 5620
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: PRM and BS oplev laser died 

The He-Ne laser which has been used for the PRM and BS oplevs were found to be dead.

According to the trend data shown below, it became dead during the dolphin issue.

(During the dolphin issue the output from the oplev QPDs are digitally zero)

oplevs.png

ELOG V3.1.3-